### v1

# Integral representations for the joint survival functions of the cumulated components of multinomial random vectors

Abstract
This paper presents a multivariate normal integral expression for the joint survival function of the cumulated components of any multinomial random vector. This result can be viewed as a multivariate analog of Equation (7) from Carter and Pollard [2], who improved Tusnády’s inequality. Our findings are based on a crucial relationship between the joint survival function of the cumulated components of any multinomial random vector and the joint cumulative distribution function of a corresponding Dirichlet distribution. We offer two distinct proofs: the first expands the logarithm of the Dirichlet density, while the second employs Laplace’s method applied to the Dirichlet integral.
keywords:
Dirichlet distribution, Gaussian approximation, Laplace’s method, multinomial distribution, multivariate normal distribution, quantile coupling
MSC:
[2020]Primary: 62E17 Secondary: 62H10, 62H12, 62E20
1 Introduction
For any integer $d\in\mathbb{N}$, the $d$-dimensional simplex and its interior are defined by
$\mathcal{S}_{d}=\big{\{}\boldsymbol{s}\in[0,1]^{d}:\|\boldsymbol{s}\|_{1}\leq 1\big{\}},\qquad\mathrm{Int}(\mathcal{S}_{d})=\big{\{}\boldsymbol{s}\in(0,1)^{d}:\|\boldsymbol{s}\|_{1}<1\big{\}},$
where $\|\boldsymbol{s}\|_{1}=\sum_{i=1}^{d}|s_{i}|$ denotes the $\ell_{1}$-norm in $\mathbb{R}^{d}$. Given a set of probability weights $\boldsymbol{p}\in\mathrm{Int}(\mathcal{S}_{d})$, the $\mathrm{Multinomial}\hskip 0.56905pt(n,\boldsymbol{p})$ probability mass function is defined, for all $\boldsymbol{k}\in\mathbb{N}_{0}^{d}\cap n\mathcal{S}_{d}$, by
$p_{n}(\boldsymbol{k})=\frac{n!}{(n-\|\boldsymbol{k}\|_{1})!\prod_{i=1}^{d}k_{i}!}\,p_{d+1}^{n-\|k\|_{1}}\prod_{i=1}^{d}p_{i}^{k_{i}},$
where $p_{d+1}=1-\|\boldsymbol{p}\|_{1}\in(0,1)$ and $n\in\mathbb{N}$. The covariance matrix of the multinomial distribution is well-known to be $n\,\Sigma_{\boldsymbol{p}}$, where $\Sigma_{\boldsymbol{p}}=\text{diag}(\boldsymbol{p})-\boldsymbol{p}\boldsymbol{p}^{\top}$; see, e.g., [7, p. 377]. From Theorem 1 of Tanabe and Sagae [9], it is also well known that $\det(\Sigma_{\boldsymbol{p}})=p_{1}\dots p_{d}\,p_{d+1}$. The centered multivariate normal density with the same covariance structure as the above multinomial distribution is defined, for all $\boldsymbol{x}\in\mathbb{R}^{d}$, by
$\phi_{\Sigma_{\boldsymbol{p}}}(\boldsymbol{x})=\frac{\exp\big{(}-\boldsymbol{x}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\,\boldsymbol{x}/2\big{)}}{\sqrt{(2\pi)^{d}\,\det(\Sigma_{\boldsymbol{p}})}}.$
Let $[d]=\{1,\ldots,d\}$. Let $I_{1}=(0,p_{1}]$ and $I_{j}=(p_{1}+\dots+p_{j-1},p_{1}+\dots+p_{j}]$ for all $j\in[d]\backslash\{1\}$. If $U_{1},\dots,U_{n}\smash{\stackrel{{\scriptstyle\mathrm{iid}}}{{\sim}}}\mathrm{Uniform}(0,1)$, then
$\boldsymbol{X}=(X_{1},\ldots,X_{d})\equiv\Big{(}\sum_{i=1}^{n}\mathds{1}_{\{U_{i}\in I_{1}\}},\ldots,\sum_{i=1}^{n}\mathds{1}_{\{U_{i}\in I_{d}\}}\Big{)}\sim\mathrm{Multinomial}\hskip 0.56905pt(n,\boldsymbol{p}).$
If $U_{(1)}\leq\dots\leq U_{(n)}$ denote the order statistics, then
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=\mathsf{P}(U_{(k_{1}+\dots+k_{i})}\leq p_{1}+\dots+p_{i},~{}\forall i\in[d])$
$\qquad=\int_{0}^{p_{1}}\int_{u_{1}}^{p_{1}+p_{2}}\dots\int_{u_{d-1}}^{p_{1}+\dots+p_{d}}n!\prod_{i=1}^{d+1}\frac{(u_{i}-u_{i-1})^{k_{i}-k_{i-1}-1}}{(k_{i}-k_{i-1}-1)!}\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{1}\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{2}\dots\hskip 1.0pt\mathrm{d}\hskip 0.75ptu_{d},$
where one defines $u_{0}=0$, $u_{d+1}=1$, $k_{0}=0$, $k_{d+1}=n+1$ in the multidimensional integral. After applying the change of variables $s_{i}=u_{i}-u_{i-1}$ and $j_{i}=k_{i}-k_{i-1}$, for all $i\in[d]$, together with $j_{d+1}=k_{d+1}-k_{d}=(n+1)-\sum_{i=1}^{d}j_{i}$, the above can be rewritten as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=\int_{0}^{p_{1}}\int_{0}^{(p_{1}-s_{1})+p_{2}}\dots\int_{0}^{\sum_{k=1}^{d-1}(p_{k}-s_{k})+p_{d}}n!\prod_{i=1}^{d+1}\frac{s_{i}^{j_{i}-1}}{(j_{i}-1)!}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
This neat equation provides a direct relationship between the joint survival function of the cumulated components of any multinomial random vector and the joint cumulative distribution function of a corresponding Dirichlet distribution.
In turn, letting $N=n-d$, $J_{i}=j_{i}-1$ for all $i\in[d+1]$, and defining the region
$\mathcal{R}_{d}=\left\{\boldsymbol{s}\in\mathcal{S}_{d}:(s_{1},s_{2},\dots,s_{i})\in\left(\sum_{k=1}^{i}p_{k}\right)\mathcal{S}_{i}~{}~{}\forall i\in[d]\right\},$ (1)
one can write
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=\int_{\mathcal{R}_{d}}\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\prod_{i=1}^{d+1}s_{i}^{J_{i}}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$ (2)
2 Quantile coupling
For any integer $m\in\mathbb{N}$, let $\lambda_{m}$ denote the error term in Stirling’s approximation for $\ln(m!)$. It is well known that, for all $m\in\mathbb{N}$,
$\ln(m!)=\frac{1}{2}\ln(2\pi m)+m\ln m-m+\lambda_{m},$ (3)
where $(12m+1)^{-1}\leq\lambda_{m}\leq(12m)^{-1}$; see, e.g., Gronwall [3].
Together with the notation introduced in Section 1, define, for all $i\in[d+1]$,
$\varepsilon_{i}=\frac{J_{i}/N-p_{i}}{p_{i}},\qquad\widetilde{\varepsilon}_{i}=p_{i}\varepsilon_{i}=J_{i}/N-p_{i},$
and also
$\Lambda_{N}$ $=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i},\qquad\Delta_{N}=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}+\Lambda_{N}-\frac{1}{2}\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (4)
$\widetilde{\gamma}(\boldsymbol{\varepsilon})$ $=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln(1+\varepsilon_{i})-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}},$ (5)
$\gamma^{\star}(\boldsymbol{s})$ $=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)-\left\{\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})\right\}.$ (6)
Theorem 1 below expresses the joint survival function of the cumulated components of any multinomial random vector in terms of a multivariate normal integral. The result can be seen as a multivariate analog of Eq. (7) of Carter and Pollard [2], who improved on Tusnády’s inequality. For an overview of the literature on Tusnády’s inequality and the most recent improvement in the bulk, see Massart [4] and Ouimet [6], respectively.
Theorem 1.
For all $\boldsymbol{k}\in\mathbb{N}_{0}^{d}\cap n\mathcal{S}_{d}$, one has
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
Useful alternative expressions for $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ and $\gamma^{\star}(\boldsymbol{s})$ can be found in (10) and (11), respectively.
Remark 1.
Theorem 1 also complements the local limit theorem for the multinomial distribution developed independently by Siotani and Fujikoshi [8] and Ouimet [5], and the related non-asymptotic approximations for Pearson’s chi square statistic by Bax and Ouimet [1].
3 Proofs
First proof of Theorem 1.
Note that $\sum_{i=1}^{d+1}J_{i}=N$. By taking the logarithm of the integrand in (2) and applying Stirling’s formula (3), one obtains
$\ln\left\{\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\prod_{i=1}^{d+1}s_{i}^{J_{i}}\right\}$ (7)
$\quad=\ln\left\{\frac{(N+d)!}{N!}\right\}+\ln(N!)-\sum_{i=1}^{d+1}\ln(J_{i}!)+\sum_{i=1}^{d+1}J_{i}\ln s_{i}$
$\quad=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}-\frac{1}{2}\ln\left\{\left(\frac{2\pi}{N}\right)^{d}\,\prod_{i=1}^{d+1}(J_{i}/N)\right\}+\sum_{i=1}^{d+1}J_{i}\ln\left(\frac{s_{i}}{J_{i}/N}\right)+\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i}$
$\quad=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}-\frac{1}{2}\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-\frac{1}{2}\ln\left\{\left(\frac{2\pi}{N}\right)^{d}\,\prod_{i=1}^{d+1}p_{i}\right\}$
$\qquad+N\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)+N\sum_{i=1}^{d+1}(J_{i}/N)\ln\left(\frac{p_{i}}{J_{i}/N}\right)+\Lambda_{N}.$
Using the Taylor expansion
$(1+x)\ln(1+x)=x+\frac{x^{2}}{2}-\frac{x^{3}}{6}+\frac{x^{4}}{12}+\mathcal{O}_{\eta}(x^{5}),\quad|x|\leq\eta<1,$
and the fact that $\widetilde{\varepsilon}_{d+1}=-\sum_{i=1}^{d}\widetilde{\varepsilon}_{i}$, one can write
$\sum_{i=1}^{d+1}(J_{i}/N)\ln\left(\frac{p_{i}}{J_{i}/N}\right)=-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (8)
where the quantity $\widetilde{\gamma}(\boldsymbol{\varepsilon})$, as defined in (5), can be expanded as follows:
$\widetilde{\gamma}(\boldsymbol{\varepsilon})$ $\equiv\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln(1+\varepsilon_{i})-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}$
$=\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}(i=j)+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}$
$\quad-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}(i=j=k)-\frac{1}{p_{d+1}^{2}}\right\}$
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}(i=j=k=\ell)+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p}}\big{(}\|\boldsymbol{\varepsilon}\|_{1}^{5}\big{)}.$
Moreover, by [9, Eq. 21], it is known that $(\Sigma_{\boldsymbol{p}}^{-1})_{ij}=p_{i}^{-1}\mathds{1}_{\{i=j\}}+p_{d+1}^{-1}$ for all $i,j\in[d]$, so
$\frac{1}{2}\sum_{i,j=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\left\{\frac{1}{p_{i}}\mathds{1}(i=j)+\frac{1}{p_{d+1}}\right\}-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}=0.$ (9)
Therefore, using the last two equations, one can rewrite $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ as follows:
$\widetilde{\gamma}(\boldsymbol{\varepsilon})$ $=-\frac{1}{6}\sum_{i,j,k=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\left\{\frac{1}{p_{i}^{2}}\mathds{1}(i=j=k)-\frac{1}{p_{d+1}^{2}}\right\}$ (10)
$\quad+\frac{1}{12}\sum_{i,j,k,\ell=1}^{d}\widetilde{\varepsilon}_{i}\widetilde{\varepsilon}_{j}\widetilde{\varepsilon}_{k}\widetilde{\varepsilon}_{\ell}\left\{\frac{1}{p_{i}^{3}}\mathds{1}(i=j=k=\ell)+\frac{1}{p_{d+1}^{3}}\right\}+\mathcal{O}_{d,\boldsymbol{p}}\big{(}\|\boldsymbol{\varepsilon}\|_{1}^{5}\big{)}.$
Now, consider $\boldsymbol{s}=(s_{1},\ldots,s_{d})\in\mathcal{R}_{d}$ as defined in (1), and let
$\delta_{i}(s)=\frac{s_{i}-p_{i}}{p_{i}},~{}~{}i\in[d],\qquad\delta_{d+1}(s)=\frac{s_{d+1}-p_{d+1}}{p_{d+1}}=-\frac{\sum_{i=1}^{d}(s_{i}-p_{i})}{p_{d+1}}.$
In particular, note that this definition yields $\sum_{i=1}^{d+1}p_{i}\delta_{i}(s_{i})=0$. Upon applying a second-order version of the Taylor expansion
$\ln(1+x)=x-\frac{x^{2}}{2}+\frac{x^{3}}{3}+\mathcal{O}_{\eta}(x^{4}),\quad|x|\leq\eta<1,$
with Lagrange remainder, there exists a vector $\boldsymbol{s}^{\star}=(s_{1}^{\star},\ldots,s_{d}^{\star})$ on the line segment joining $\boldsymbol{s}$ and $\boldsymbol{p}=(p_{1},\ldots,p_{d})$ such that
$\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\{1+\delta_{i}(s_{i})\}$
$\quad=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\delta_{i}(s_{i})-\frac{1}{2}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}$
$\quad=\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\delta_{i}(s_{i})-\frac{1}{2}\sum_{i=1}^{d+1}p_{i}\{\delta_{i}(s_{i})\}^{2}-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}$
$\quad=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}},$
where $\widetilde{\boldsymbol{\varepsilon}}=(\widetilde{\varepsilon}_{1},\ldots,\widetilde{\varepsilon}_{d})$. Using the last equation, the quantity $\gamma^{\star}(\boldsymbol{s})$, as defined in (6), can be rewritten as
$\gamma^{\star}(\boldsymbol{s})$ $\equiv\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)-\left\{\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})\right\}$ (11)
$=-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}.$
It readily follows that
$\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)$ $=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})+\gamma^{\star}(\boldsymbol{s})$ (12)
$=\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{J}/N)^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{J}/N)+\gamma^{\star}(\boldsymbol{s}).$
Therefore, by putting (8) and (12) back into (7), and exponentiating, one gets
$\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\prod_{i=1}^{d+1}s_{i}^{J_{i}}$ $=\exp\left\{\Delta_{N}+N\gamma^{\star}(\boldsymbol{s})\right\}\frac{\exp\big{\{}-N(\boldsymbol{s}-\boldsymbol{J}/N)^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{J}/N)/2\big{\}}}{\sqrt{(2\pi/N)^{d}\det(\Sigma_{\boldsymbol{p}})}}$
$=\exp\left\{\Delta_{N}+N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{J}/N-\boldsymbol{s})\}$
$=\exp\left\{\Delta_{N}+N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\},$
where recall $\Delta_{N}=\ln\{(N+d)!/(N!N^{d})\}+\Lambda_{N}-(1/2)\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon})$ from (4). Putting this last equation in (2) yields
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left[N\gamma^{\star}(\boldsymbol{s})\right]N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
This concludes the proof. ∎
Second proof of Theorem 1.
For all $\boldsymbol{s}\in\mathcal{S}_{d}$, define
$H(\boldsymbol{s})=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln s_{i}=\sum_{i=1}^{d+1}\frac{J_{i}}{N}\ln s_{i}.$
For all $\boldsymbol{k}\in\mathbb{N}_{0}^{d}\cap n\mathcal{S}_{d}$, the integral representation of the joint survival function of the cumulated components of any multinomial random vector derived in (2) can be rewritten as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])=\frac{(N+d)!}{N!}\times\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}\int_{\mathcal{R}_{d}}\exp\left\{NH(\boldsymbol{s})\right\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$ (13)
By Stirling’s formula (3), note that
$m!=\sqrt{2\pi m}\,\exp(m\ln m-m+\lambda_{m}),$
where $(12m+1)^{-1}\leq\lambda_{m}\leq(12m)^{-1}$. Since $\sum_{i=1}^{d+1}J_{i}=N$ and $\Lambda_{N}=\lambda_{N}-\sum_{i=1}^{d+1}\lambda_{i}$, one deduces that
$\frac{N!}{\prod_{i=1}^{d+1}J_{i}!}=\frac{\exp(\Lambda_{N})}{\sqrt{(2\pi N)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\exp\left\{-NH(\boldsymbol{J}/N)\right\},$
Therefore, one can rewrite (13) as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$ (14)
$\qquad=\frac{(N+d)!}{N!N^{d}}\times\frac{N^{d/2}\exp(\Lambda_{N})}{\sqrt{(2\pi)^{d}\prod_{i=1}^{d+1}\{p_{i}(1+\varepsilon_{i})\}}}\int_{\mathcal{R}_{d}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{J}/N)\right\}\right]\,\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
Decompose the integral above as
[…]
Using the quantity $\widetilde{\gamma}(\boldsymbol{\varepsilon})$ as defined in (5), one has
$H(\boldsymbol{p})-H(\boldsymbol{J}/N)=-\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln(1+\varepsilon_{i})=-\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\widetilde{\gamma}(\boldsymbol{\varepsilon}),$ (16)
where recall $\widetilde{\boldsymbol{\varepsilon}}=(p_{1}\varepsilon_{1},\ldots,p_{d}\varepsilon_{d})^{\top}$, $\widetilde{\varepsilon}_{d+1}=p_{d+1}\varepsilon_{d+1}$, and $\Sigma_{\boldsymbol{p}}^{-1}=\{p_{i}^{-1}\mathds{1}(i=j)+p_{d+1}^{-1}\}_{1\leq i,j\leq d}$. Recall from (4) that
$\Delta_{N}=\ln\left\{\frac{(N+d)!}{N!N^{d}}\right\}+\Lambda_{N}-\frac{1}{2}\sum_{i=1}^{d+1}\ln(1+\varepsilon_{i})-N\widetilde{\gamma}(\boldsymbol{\varepsilon}),$
so one can rewrite (14) as
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=\frac{N^{d/2}\exp(\Delta_{N})}{\sqrt{(2\pi)^{d}\det(\Sigma_{\boldsymbol{p}})}}\int_{\mathcal{R}_{d}}\exp\left[N\left\{H(\boldsymbol{s})-H(\boldsymbol{p})\right\}-\frac{N}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}\right]\,\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
Using (11) and (12), one has
$H(\boldsymbol{s})-H(\boldsymbol{p})$ $=\sum_{i=1}^{d+1}\frac{J_{i}}{N}\ln(s_{i})-\sum_{i=1}^{d+1}\frac{J_{i}}{N}\ln(p_{i})=\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\ln\left(\frac{s_{i}}{p_{i}}\right)$
$=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})$
$\qquad-\frac{1}{2}\sum_{i=1}^{d+1}\widetilde{\varepsilon}_{i}\{\delta_{i}(s_{i})\}^{2}+\frac{1}{3}\sum_{i=1}^{d+1}p_{i}(1+\varepsilon_{i})\frac{\{\delta_{i}(s_{i})\}^{3}}{\{1+\delta_{i}(s_{i}^{\star})\}^{3}}$
$=\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{p})^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{p})+\gamma^{\star}(\boldsymbol{s})$
$=\frac{1}{2}\widetilde{\boldsymbol{\varepsilon}}^{\top}\Sigma_{\boldsymbol{p}}^{-1}\widetilde{\boldsymbol{\varepsilon}}-\frac{1}{2}(\boldsymbol{s}-\boldsymbol{J}/N)^{\top}\Sigma_{\boldsymbol{p}}^{-1}(\boldsymbol{s}-\boldsymbol{J}/N)+\gamma^{\star}(\boldsymbol{s}).$
Putting this in the previous equation yields
$\mathsf{P}(X_{1}+\dots+X_{i}\geq k_{1}+\dots+k_{i},~{}\forall i\in[d])$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{J}/N-\boldsymbol{s})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}$
$\qquad=e^{\Delta_{N}}\int_{\mathcal{R}_{d}}\exp\left\{N\gamma^{\star}(\boldsymbol{s})\right\}N^{d/2}\phi_{\Sigma_{\boldsymbol{p}}}\{N^{1/2}(\boldsymbol{p}-\boldsymbol{s}+\widetilde{\boldsymbol{\varepsilon}})\}\hskip 1.0pt\mathrm{d}\hskip 0.75pt\boldsymbol{s}.$
This concludes the proof. ∎
Remark 2.
The choice of decomposition in (15) can be explained as follows. The Hessian matrix of $H$ at $\boldsymbol{s}$ is equal to
$\left(-\frac{J_{i}}{Ns_{i}^{2}}\mathds{1}(i=j)-\frac{J_{d+1}}{Ns_{d+1}^{2}}\right)_{1\leq i,j\leq d},$
which is (symmetric) negative definite, so that $H$ is strictly concave on $\mathrm{Int}(\mathcal{S}_{d})$. Given the strict concavity and the fact that $\overline{\boldsymbol{s}}=\boldsymbol{J}/N\in\mathrm{Int}(\mathcal{S}_{d})$ is the unique point satisfying
$\nabla H(\overline{\boldsymbol{s}})=\left(\frac{J_{i}}{N\,\overline{s}_{i}}-\frac{J_{d+1}}{N\,\overline{s}_{d+1}}\right)_{1\leq i\leq d}=\boldsymbol{0}$
in the interior of the simplex, the global maximum of $H$ on $\mathrm{Int}(\mathcal{S}_{d})$ is achieved at $\boldsymbol{J}/N$. When the point $\boldsymbol{k}\in\mathbb{N}_{0}\cap n\mathrm{Int}(\mathcal{S}_{d})$ is such that $\boldsymbol{J}/N$ falls outside the range of integration $\mathcal{R}_{d}$ defined in (1), then $H$ maximizes at $\boldsymbol{s}=\boldsymbol{p}$ inside the region of integration $\mathcal{R}_{d}$. The decomposition is therefore equivalent to an application of Laplace’s method provided some restrictions on $\boldsymbol{k}$ are met.
Funding
F. Ouimet’s current postdoctoral fellowship is funded through the Canada Research Chairs Program (Grant 950-231937 to Christian Genest) and the Natural Sciences and Engineering Research Council of Canada (Grant RGPIN-2024-04088 to Christian Genest).